# Semiregular abstract polyhedra with trivial facet stabilizer

### v1

@@ 6 Constructing polyhedra with symmetry type graph in family 1
    Let $i$ and $j$ be such that $\Psi_{i}=\Psi_{j}$ with $10\leq i<j\leq n\ell-$ 1. Then $\Psi_{j}=\prod_{k=i}^{j-1}r_{k}\Psi_{i}$. Since $\Psi_{i}$ and $\Psi_{j}$ have the same vertex, we get that $\left\lfloor(i-1)/n\right\rfloor=\left\lfloor(j-1)/n\right\rfloor$.In particular $j-i<n$, so $r_{i}$ appears exactly once in the product $\prod_{k=i}^{j-1}r_{k}$. But since $\mathcal{P}$ is a polytope, this implies that $\prod_{k=i}^{j-1}r_{k}$ must change the $i-face$ of every flag, so we have reached a contradiction.∎
Going back to the case where $\mathcal{P}$ is a regular $n$-polytope with simple Petrie polygons, using that $\Phi\rho_{i}=r_{i}\Phi$ and the fact that auromorphisms commute with monodromies one can see that
@@
So the sequence $S$ in 6.2 is the same as $\{\Phi\prod_{i=0}^{k-1}\rho_{i}\}_{k=0}^{n\ell-1}$ where the subscripts are taken modulo $n$. Applying the lema we get that $\Phi\prod_{i=0}^{k-1}\rho_{i}=\Phi$ if and only if $k$ is divisible by $n\ell$. In particular, if $\prod_{i=0}^{k-1}\rho_{i}=1$ then $k$ is divisible by $n$.
Now let $W$ be a path with darts of colors 1 and 2. Without loss of generality, we may assume that $W$ is reduced, that is, that it does not have any pair of consecutive inverse darts. This is equivalent to saying that $W$ alternates between colors 1 and 2. Let $W^{\prime}$ be the path that follows the same color sequence as $W$ but starting at the base flag of $\mathcal{X}$. Notice that the voltage of $W$ is a conjugate of the voltage of $W^{\prime}$, so $\xi(W)=1$ if and only if $\xi(W^{\prime})=1$. But we have proved that $\xi(W^{\prime})=1$ if and only if $W^{\prime}$ ends at either the base flag or its 1-adjacent flag, and since $\mathcal{X}$ is regular, the same must be true for $W$.
@@
Let $\mathcal{P}$ be a regular $n$-polytope with simple Petrie polygons. Suppose that $\mathcal{P}$ is non-degenerate and non-self-dual. Let $(\mathcal{X},\Gamma,\xi)$ be the voltage premaniplex where $\mathcal{X}$ is an the premaniplex of family 1 with $n$-sides, $\Gamma$ is the automorphism group of $\mathcal{P}$ with distinguished generators $\{\rho_{0},\ldots,\rho_{n-1}\}$, and $\xi$ is the voltage assignment that assigns voltage $\rho_{i}$ to the darts starting at the $i$-th 1-face (as in Figure 6). Then $\mathcal{X}^{\xi}$ is a polyhedron with symmetry type graph isomorphic to $\mathcal{X}$. In particular $\mathcal{X}^{\xi}$ is vertex-transitive, facet-transitive and it has trivial facet stabiliizer.
@@ 7 Constructing polytopes with symmetry type graph in family 2
To construct polytopes with symmetry type graphs in family 2 we procede in a similar fashion. Let $\mathcal{X}$ be the premaniplex in family 2 constructed from a $4k$-gon (see Figure 5). We enumerate the 1-faces of $\mathcal{X}$ cyclically starting with $e_{0}$ as the 1-face of the base flag. Note that an edge of color 2 joins a flag in the 1-face $e_{i}$ with one in the edge $e_{2k+i}$ (with the suscripts taken modulo $4k$). We assign trivial voltage to the darts of colors 0 and 1 to ensure that the facets of the derived graph are $4k$-gons. If we assign a voltage $\alpha_{i}$ to a dart going from edge $e_{i}$ to edge $e_{2k+i}$, we must assign the same voltage to the other dart that goes from $e_{i}$ to $e_{2k+i}$. This is to ensure that the derived graph is a maniplex.
To assign the voltages to the darts of color 2, this time we use the automorphism group of a chiral polytope. A maniplex is chiral if it has 2 orbits on flags and adjacent flags are in different orbits. In 2010 Pellicer [15] proved that there are abstract chiral polytopes in every rank greater or equal to 3. Let $\mathcal{P}$ be a chiral polytope of rank $12k+$ 1 and let $\Phi$ be a base flag of $\mathcal{P}$. Because of chirality, for every $i$ in $\{1,2,\ldots,2k\}$, there is an automorphism $\tau_{i}$ that maps $\Phi$ to $r_{i}r_{0}\Phi$. It is worth noting that $\tau_{i}$ is an involution if and only if $i\neq 1$. For $i\in\{0,1,\ldots,2k-1\}$ we will assign the voltage $\tau_{i+1}$ to the darts of color 2 in $\mathcal{X}$ that connect the 1-face $e_{i}$ to $e_{2k+i}$. This voltage assignment is ilustrated in Figure 7, where once again, the labels of the edges of the polygon correspond to the voltages of the darts of color 2 that start on that edge.
@@
Using the fact that $\tau_{1}$ is the only $\tau_{i}$ that is not an involution we can see that the only possible automorphisms of $\mathcal{X}$ that could induce an automorphism of $\Gamma(\mathcal{P})$ are the identiy and the half turn. However, it is known (see [15, Proposition 2.2], for example) that if $\mathcal{P}$ is a chiral polytope, there is no automorphism of $\Gamma(\mathcal{P})$ that maps $\tau_{1}$ to $\tau_{1}^{-1}$ and fixes every other $\tau_{i}$. Therefore, using 4.1 we get that no automorphism of $\mathcal{X}$ lifts, and therefore $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$.
If we look at a path $W$ of alternating colors 2 and 1 that starts on the left flag of $e_{0}$ and uses $j$ edges of color 2, we will see that its voltage is $\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}$, where the subscript of $\tau_{i}$ is taken modulo $2k$ (using $\tau_{2k}$ when $i\equiv 0\mod 2k$) and $\epsilon_{i}$ is 1 unless $i\equiv 2k+1\mod 4k$, in which case it is $-1$. The path $W$ ends in $\Phi$ or $\Phi^{1}$ whenever it uses a multiple of $4k$ edges of color 2. So using 6.1, we get that in order to find examples where the derived graph is polytopal, we need to make sure that $\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}=1$ implies that $j$ is divisible by $4k$.
@@
Let $\mathcal{P}$ be a chiral $(2k+1)$-polytope with a base flag $\Phi$, and let $\tau_{i}$ be such that $\Phi\tau_{i}=r_{i}r_{0}\Phi$ for $1\leq i\leq 2k$ and let $\epsilon_{i}$ be defined as above. Let $j$ be a positive integer. If $j\not\equiv\left\lfloor(j-1)/(2k)\right\rfloor\mod 2$, then
@@
where the subscript of $r_{i}$ is taken modulo $12k+$ 1 and the subscript of $\tau_{i}$ modulo $2k$. On the other hand, if $j\equiv\left\lfloor(j-1)/(2k)\right\rfloor\mod 2$, then
@@
  • Proof
    We will prove this in a ‘‘hand-wavy’’ manner. The reader may formalize this proof using induction and making several cases. For $j=1$ we get $\Phi\tau_{1}=r_{1}r_{0}\Phi$ by definition of $\tau_{1}$. Now suppose the result is true for $j-1$. Using the definition of $\tau_{j}$ and the fact that automorphisms commute with monodromies we get
@@
    Recall that the subscripts of $\tau_{i}$ and $\tau_{j}$ are taken modulo $2k$, and therefore, the subscript $j$ of $r_{j}$ is also taken modulo $2k$. If $j\not\equiv 1\mod 2k$, we see that the exponent $\epsilon_{j}$ is irrelevant and the $r_{0}$ gets canceled or not depending on wether $\Phi\prod_{i=1}^{j-1}\tau_{i}^{\epsilon_{i}}$ has an $r_{0}$ on the left or not. Notice that for $j<2k+1$ there is alternation between these two cases, where the even terms are the ones with an $r_{0}$ on the left. In particular $\Phi\prod_{i=1}^{2k}\tau_{i}^{\epsilon_{i}}=r_{0}\prod_{i=0}^{2k}r_{i}\Phi$. Then for $j=2k+1$ we get
@@
    Notice how at this point the alternation pauses momentarily and we get two consecutive products with an $r_{0}$ on the left. Notice also that the superscript of the product increases by 2 rather than 1. After this point alternation resumes, but now the odd terms are the ones with an $r_{0}$ on the left. In particular $\Phi\prod_{i=1}^{4k}\tau_{i}^{\epsilon_{i}}=\prod_{i=0}^{4k+1}r_{i}\Phi$. Now for $j=4k+1$ we get:
    $\Phi\prod_{i=1}^{4k+1}\tau_{i}^{\epsilon_{i}}$ $=$ $(r_{1}r_{0})^{\epsilon_{4k+1}}(\Phi\prod_{i=1}^{4k+1}\tau_{i}^{\epsilon_{i}})$
@@
Let $\mathcal{P}$ be a chiral polytope of rank $12k+$ 1 with a base flag $\Phi$ with simple Petrie polygons. Similarily to the case for family 1, we take a path $W$ in $\mathcal{X}$ starting from the base flag, alternating darts of color 2 and 1, starting with color 1 and using a total of $j$ darts of color 2. Using 7.1 we see that the voltage $W$ acts as a $2t$-step rotation in the Petrie polygon $P$ whenever $W$ uses $4kt$ edges of color 2.
@@
If $j\not\equiv\left\lfloor(j-1)/(2k)\right\rfloor\mod 2$, then $\Phi\xi(W)=\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi$. Let $t=j+\left\lfloor(j-1)/(2k)\right\rfloor+1$. We can use the same argument as we used for family 1 to prove that $\Phi=\prod_{i=0}^{t-1}r_{i}\Phi$ implies that $t$ is divisible by the rank, which in this case is $12k+$ 1. A simple substitution shows that when $j=2km$ for some positive $m$, then $t=(2k+1)m$. Since $t$ is defined by a monotonous function on $j$, the converse is also true. This shows that if $\prod_{i=0}^{t-1}r_{i}$ fixes the base flag, then $j$ is divisible by $2k$. Since $j\not\equiv\left\lfloor(j-1)/2k\right\rfloor\mod 2$, then $j/(2k)\equiv j\equiv 0\mod 2$, and therefore $j$ must be a multiple of $4k$.
Now suppose $j\equiv\left\lfloor(j-1)/2\right\rfloor\mod 2$ and $\Phi=\Phi\xi(W)=r_{0}\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi$. Then $r_{0}\Phi=\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi$. Let $t=j+\left\lfloor(j-1)/(2k)\right\rfloor+1$ and let $t^{\prime}$ be its residue modulo $(2k+1)\ell$, where $\ell$ is the length of $P$. Then, by definition of $\ell$ we see that $r_{0}\Phi=\prod_{i=0}^{t-1}\Phi=\prod_{i=0}^{t^{\prime}-1}r_{i}\Phi$. 6.2 tells us that $t^{\prime}=1$, in particular $t\equiv 1\mod 2k+1$. We have established that $t$ is divisible by $12k+$ 1 if and only if $j$ is divisible by $2k$. If $j\equiv r\mod 2k$, for $1\leq r<2k$, a substitution shows that $t\equiv r+1\mod 2k+1$. Therefore, it is not possible that $t\equiv 1\mod 2k+1$. This proves that this case cannot happen.
@@
Let $\mathcal{P}$ be a chiral $n$-polytope with simple Petrie polygons. Suppose that $\mathcal{P}$ is non-degenerate and non-self-dual. Let $(\mathcal{X},\Gamma,\xi)$ be the voltage premaniplex where $\mathcal{X}$ is the premaniplex of family 2 with $4k$ sides, $\Gamma$ is the automorphism group of $\mathcal{P}$ with distinguished generators $\{\tau_{1},\ldots,\tau_{2k}\}$, and $\xi$ is the voltage assignment that assigns voltage $\tau_{i+1}$ to the darts starting at the $i$-th 1-face for $i<2k$ (as in Figure 7). Then $\mathcal{X}^{\xi}$ is a polyhedron with symmetry type graph isomorphic to $\mathcal{X}$. In particular $\mathcal{X}^{\xi}$ is vertex-transitive, facet-transitive and it has trivial facet stabiliizer.
8 Interpreting with voltage operations

### v2

@@ 6 Constructing polyhedra with symmetry type graph in family 1
Let $i$ and $j$ be such that $\Psi_{i}=\Psi_{j}$ with $0\leq i<j\leq n\ell-1$. Then $\Psi_{j}=\prod_{k=i}^{j-1}r_{k}\Psi_{i}$. Since $\Psi_{i}$ and $\Psi_{j}$ have the same vertex, we get that $\left\lfloor(i-1)/n\right\rfloor=\left\lfloor(j-1)/n\right\rfloor$. In particular ji<nj-i<n, so $r_{i}$ appears exactly once in the product $\prod_{k=i}^{j-1}r_{k}$. But since $\mathcal{P}$ is a polytope, this implies that $\prod_{k=i}^{j-1}r_{k}$ must change the $i$-face of every flag, so we have reached a contradiction.∎
Going back to the case where $\mathcal{P}$ is a regular $n$-polytope with simple Petrie polygons, using that $\Phi\rho_{i}=r_{i}\Phi$ and the fact that automorphisms commute with monodromies one can see that
@@
So the sequence $S$ in ˜6.2 is the same as $\{\Phi\prod_{i=0}^{k-1}\rho_{i}\}_{k=0}^{n\ell-1}$ where the subscripts are taken modulo $n$. Applying the lemma we get that $\Phi\prod_{i=0}^{k-1}\rho_{i}=\Phi$ if and only if $k$ is divisible by $n\ell$. In particular, if $\prod_{i=0}^{k-1}\rho_{i}=1$ then $k$ is divisible by $n$.
Now let $W$ be a path with darts of colors 1 and 2. Without loss of generality, we may assume that $W$ is reduced, that is, that it does not have any pair of consecutive inverse darts. This is equivalent to saying that $W$ alternates between colors 1 and 2. Let $W^{\prime}$ be the path that follows the same color sequence as $W$ but starting at the base flag of $\mathcal{X}$. Notice that the voltage of $W^{\prime}$ is a conjugate of either the voltage of $W$ or the voltage of $W^{-1}$ (depending on whether $W$ starts at a left or a right flag), so $\xi(W)=1$ if and only if $\xi(W^{\prime})=1$. But we have proved that $\xi(W^{\prime})=1$ if and only if $W^{\prime}$ ends at either the base flag or its 1-adjacent flag, and since $\mathcal{X}$ is regular, the same must be true for $W$.
@@
Let $\mathcal{P}$ be a regular $n$-polytope with simple Petrie polygons. Suppose that $\mathcal{P}$ is non-degenerate and non-self-dual. Let $(\mathcal{X},\Gamma,\xi)$ be the voltage premaniplex where $\mathcal{X}$ is the premaniplex of family 1 with $n$ sides, $\Gamma$ is the automorphism group of $\mathcal{P}$ with distinguished generators $\{\rho_{0},\ldots,\rho_{n-1}\}$, and $\xi$ is the voltage assignment that assigns voltage $\rho_{i}$ to the darts starting at the $i$-th 1-face (as in Figure˜6). Then $\mathcal{X}^{\xi}$ is a polyhedron with symmetry type graph isomorphic to $\mathcal{X}$. In particular $\mathcal{X}^{\xi}$ is vertex-transitive, facet-transitive and it has trivial facet stabilizer.
Example 6.4.
Let $\mathcal{P}$ be a 3-dimensional cube, and let $\rho_{0}$, $\rho_{1}$ and $\rho_{2}$ be the distinguished generators for its automorphism group. Let $\mathcal{X}$ be the premaniplex in family 1 with 3 sides. Then, the derived graph from the voltage graph in Figure˜6 is a semiregular facet-transitive polyhedron with 48 triangular faces, 72 edges and 8 vertices of degree 18. It has genus 9 and it has trivial facet stabilizer. If we replace the cube with the octahedron we get the same resulting polyhedron.
Analogously, both the icosahedron and the dodecahedron give us a polyhedron with 120 triangular faces, 180 edges and 12 vertices of degree 30. It has genus 25 and trivial facet stabilizer.
Finally, if $\mathcal{P}$ is the tetrahedron the result is a polyhedron with 24 triangular faces, 36 edges and 6 vertices of degree 12. It has genus 4 and its facet stabilizer has order two, as the tetrahedron is self-dual.
Notice that the degree of the vertices in $\mathcal{O}(\mathcal{P})$ is $n$ times the length of the Petrie polygons of $\mathcal{P}$.
@@ 7 Constructing polytopes with symmetry type graph in family 2
To construct polytopes with symmetry type graphs in family 2 we proceed in a similar fashion. Let $\mathcal{X}$ be the premaniplex in family 2 constructed from a $4k$-gon (see Figure˜5). Once again, we enumerate the 1-faces of $\mathcal{X}$ so that $e_{i}$ is the 1-face of the flag $(r_{1}r_{0})^{i}\Phi$ where $\Phi$ is the base flag. Note that an edge of color 2 joins a flag in the 1-face $e_{i}$ with one in the edge $e_{2k+i}$ (with the subscripts taken modulo $4k$). We assign trivial voltage to the darts of colors 0 and 1 to ensure that the facets of the derived graph are $4k$-gons. If we assign a voltage $\alpha_{i}$ to a dart going from edge $e_{i}$ to edge $e_{2k+i}$, we must assign the same voltage to the other dart that goes from $e_{i}$ to $e_{2k+i}$. This is to ensure that the derived graph is a maniplex.
To assign the voltages to the darts of color 2, this time we use the automorphism group of a chiral polytope. A maniplex is chiral if it has 2 orbits on flags and adjacent flags are in different orbits. In 2010 Pellicer [19] proved that there are abstract chiral polytopes in every rank greater or equal to 3. Let $\mathcal{P}$ be a chiral polytope of rank $2k+1$ and let $\Phi$ be a base flag of $\mathcal{P}$. Because of chirality, for every $i$ in $\{1,2,\ldots,2k\}$, there is an automorphism $\tau_{i}$ that maps $\Phi$ to $r_{i}r_{0}\Phi$. It is worth noting that $\tau_{i}$ is an involution if and only if $i\neq 1$ and $\tau_{i}\tau_{j}$ is an involution if and only if |ij|>1|i-j|>1. This is a consequence of the known fact that there are no degenerate chiral polytopes [20, Lemma 2.3.16]. For $i\in\{0,1,\ldots,2k-1\}$ we will assign the voltage $\tau_{i+1}$ to the darts of color 2 in $\mathcal{X}$ that connect the 1-face $e_{i}$ to $e_{2k+i}$. This voltage assignment is illustrated in Figure˜7, where once again, the labels of the edges of the polygon correspond to the voltages of the darts of color 2 that start on that edge.
@@
Using the fact that $\tau_{1}$ is the only $\tau_{i}$ that is not an involution we can see that the only possible automorphisms of $\mathcal{X}$ that could induce an automorphism of $\Gamma(\mathcal{P})$ are the identity, the half turn, the reflection that fixes $e_{0}$ and $e_{2k}$, and the reflection that fixes $e_{k}$ and $e_{3k}$.
It is known (see [18, Corollary 2.5]) that if $\mathcal{P}$ is a chiral polytope, there is no automorphism of $\Gamma(\mathcal{P})$ that maps $\tau_{1}$ to $\tau_{1}^{-1}$ and fixes every other $\tau_{i}$. Therefore, using ˜4.1 we get that the half-turn does not lift.
The reflection that fixes $e_{0}$ fixes edges with voltage $\tau_{1}$ and for $2\leq i\leq 2k$, it swaps edges with voltage $\tau_{i}$ with edges with voltage $\tau_{2k+2-i}$. Notice that $\tau_{1}\tau_{2}$ is not an involution but $\tau_{1}\tau_{2k}$ is whenever $k\geq 2$. Therefore, the reflection that fixes $e_{0}$ does not lift for $k\geq 2$. Notice that for $k=1$, the reflection that fixes $e_{0}$ induces the identity automorphism in $\Gamma(\mathcal{P})$, so in that case it does lift and the derived maniplex has non-trivial facet stabilizer (the stabilizer would be the group of order of order 2 generated by the lift of this reflection).
The reflection that fixes $e_{k}$ swaps edges with voltage $\tau_{1}$ with edges with voltage $\tau_{1}^{-1}$, and it swaps edges with voltage $\tau_{i}$ with edges with voltage $\tau_{2k+2-i}$. For $k=1$, this reflection cannot induce an automorphism of $\Gamma(\mathcal{P})$, since it would fix $\tau_{2}$ while inverting $\tau_{1}$. For $k\geq 2$, we see that $\tau_{1}\tau_{2}$ is not an involution while $\tau_{1}^{-1}\tau_{2k}$ is, so once again this reflection does not lift.
We have proved that for $k\geq 2$ no non-trivial automorphism of $\mathcal{X}$ lifts, and therefore $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$.
Now we want to see when our examples are polytopes.
If we look at a path $W$ of alternating colors 2 and 1 (starting with color 2) that starts on the left flag of $e_{0}$ and uses $j$ edges of color 2, we will see that its voltage is $\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}$, where the subscript of $\tau_{i}$ is taken modulo $2k$ (using $\tau_{2k}$ when $i\equiv 0\mod 2k$) and
[…]
The path $W$ ends in $\Phi$ or $\Phi^{1}$ whenever it uses a multiple of $4k$ edges of color 2. So using ˜6.1, we get that in order to find examples where the derived graph is polytopal, we need to make sure that $\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}=1$ implies that $j$ is divisible by $4k$.
@@
Let $\mathcal{P}$ be a chiral $(2k+1)$-polytope with base flag $\Phi$, and let $\tau_{i}$ be such that $\Phi\tau_{i}=r_{i}r_{0}\Phi$ for $1\leq i\leq 2k$ and let $\epsilon_{i}$ be defined as in Equation˜1. Let $j$ be a positive integer. If $j\not\equiv\left\lfloor(j-1)/(2k)\right\rfloor\mod 2$, then
@@
where the subscript of $r_{i}$ is taken modulo $2k+1$ and the subscript of $\tau_{i}$ modulo $2k$. On the other hand, if $j\equiv\left\lfloor(j-1)/(2k)\right\rfloor\mod 2$, then
@@
We will prove this in a ‘‘hand-wavy’’ manner. The reader may formalize this proof using induction and making several cases. For $j=1$ we get $\Phi\tau_{1}=r_{1}r_{0}\Phi$ by definition of $\tau_{1}$. Now suppose the result is true for j1j-1. Using the definition of $\tau_{j}$ and the fact that automorphisms commute with monodromies we get
@@
Recall that the subscripts of $\tau_{i}$ and $\tau_{j}$ are taken modulo $2k$, and therefore, the subscript $j$ of $r_{j}$ is also taken modulo $2k$. If $j\not\equiv 1\mod 2k$, we see that the exponent $\epsilon_{j}$ is irrelevant and the $r_{0}$ gets canceled or not depending on whether $\Phi\prod_{i=1}^{j-1}\tau_{i}^{\epsilon_{i}}$ has the factor $r_{0}$ on the left or not. Notice that for $j<2k+1$ there is alternation between these two cases, where the even terms are the ones with an $r_{0}$ on the left. In particular $\Phi\prod_{i=1}^{2k}\tau_{i}^{\epsilon_{i}}=r_{0}\prod_{i=0}^{2k}r_{i}\Phi$. Then for $j=2k+1$ we get
@@
Notice how at this point the alternation pauses momentarily and we get two consecutive products with an $r_{0}$ on the left. Notice also that the superscript of the product of $r_{i}$ increases by 2 rather than 1. After this point alternation resumes, but now the odd terms are the ones with an $r_{0}$ on the left. In particular $\Phi\prod_{i=1}^{4k}\tau_{i}^{\epsilon_{i}}=\prod_{i=0}^{4k+1}r_{i}\Phi$. Now for $j=4k+1$ we get:
$\Phi\prod_{i=1}^{4k+1}\tau_{i}^{\epsilon_{i}}$ $=$ $(r_{1}r_{0})^{\epsilon_{4k+1}}(\Phi\prod_{i=1}^{4k}\tau_{i}^{\epsilon_{i}})$
@@
Let $\mathcal{P}$ be a chiral polytope of rank $2k+1$ with base flag $\Phi$ with simple Petrie polygons. Similarly to the case for family 1, we take a path $W$ in $\mathcal{X}$ starting from the base flag, alternating darts of color 2 and 1, starting with color 1 and using a total of $j$ darts of color 2. Using ˜7.1 we see that the voltage $W$ acts as a $2t$-step rotation in the Petrie polygon $P$ whenever $W$ uses $4kt$ edges of color 2.
@@
If $j\not\equiv\left\lfloor(j-1)/(2k)\right\rfloor\mod 2$, then $\Phi\xi(W)=\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi$. Let $t=j+\left\lfloor(j-1)/(2k)\right\rfloor+1$. We can use the same argument as we used for family 1 to prove that $\Phi=\prod_{i=0}^{t-1}r_{i}\Phi$ implies that $t$ is divisible by the rank, which in this case is $2k+1$. A simple substitution shows that when $j=2km$ for some positive $m$, then $t=(2k+1)m$. Since $t$ is defined by a monotone function on $j$, the converse is also true. This shows that if $\prod_{i=0}^{t-1}r_{i}$ fixes the base flag, then $j$ is divisible by $2k$. Since $j\not\equiv\left\lfloor(j-1)/2k\right\rfloor\mod 2$, then $j/(2k)\equiv j\equiv 0\mod 2$, and therefore $j$ must be a multiple of $4k$.
Now suppose $j\equiv\left\lfloor(j-1)/2\right\rfloor\mod 2$ and $\Phi=\Phi\xi(W)=r_{0}\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi$. Then $r_{0}\Phi=\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi$. Let $t=j+\left\lfloor(j-1)/(2k)\right\rfloor+1$ and let $t^{\prime}$ be its residue modulo $(2k+1)\ell$, where $\ell$ is the length of $P$. Then, by definition of $\ell$ we see that $r_{0}\Phi=\prod_{i=0}^{t-1}r_{i}\Phi=\prod_{i=0}^{t^{\prime}-1}r_{i}\Phi$. ˜6.2 tells us that $t^{\prime}=1$, in particular $t\equiv 1\mod 2k+1$. We established earlier that $t$ is divisible by $2k+1$ if and only if $j$ is divisible by $2k$. If $j\equiv r\mod 2k$, for $1\leq r<2k$, a substitution shows that $t\equiv r+1\mod 2k+1$. Therefore, it is not possible that $t\equiv 1\mod 2k+1$. This proves that this case cannot happen.
@@
Let $k\geq 2$ and let $\mathcal{P}$ be a chiral $(2k+1)$-polytope with simple Petrie polygons. Let $(\mathcal{X},\Gamma,\xi)$ be the voltage premaniplex where $\mathcal{X}$ is the premaniplex of family 2 with $4k$ sides, $\Gamma$ is the automorphism group of $\mathcal{P}$ with distinguished generators $\{\tau_{1},\ldots,\tau_{2k}\}$, and $\xi$ is the voltage assignment that assigns voltage $\tau_{i+1}$ to the darts starting at the $i$-th 1-face for $i<2k$ (as in Figure˜7). Then $\mathcal{X}^{\xi}$ is a polyhedron with symmetry type graph isomorphic to $\mathcal{X}$. In particular $\mathcal{X}^{\xi}$ is vertex-transitive, facet-transitive and it has trivial facet stabilizer.
Example 7.3.
Let $G$ be the group with presentation
$\langle\sigma_{1},\sigma_{2},\sigma_{3},\sigma_{4}$ $|\mathcal{R}(5),\sigma_{1}^{12}=\sigma_{2}^{6}=\sigma_{3}^{6}=\sigma_{4}^{3}=[\sigma_{1},\sigma_{2}^{2}]=[\sigma_{4},\sigma_{3}^{2}]=(\sigma_{2}^{-1}\sigma_{3})^{4}$
$=\sigma_{1}^{-1}\sigma_{2}\sigma_{3}^{3}\sigma_{1}^{-1}\sigma_{3}^{2}\sigma_{2}\sigma_{3}^{-1}=1\rangle$
where $\mathcal{R}(5)$ is a set of relations satisfied by every chiral 5-maniplex (see [1, Section 4]). In [1, Example 4.20] it is shown that $G$ is the automorphism group of a chiral 5-polytope $\mathcal{P}$ with 41472 flags. The generator $\sigma_{i}$ maps the base flag $\Phi$ to $\Phi^{i,i-1}$, and one can show that if $\tau_{i}$ is defined as the automorphism of $\mathcal{P}$ that maps $\Phi$ to $\Phi^{0,i}$ then $\sigma_{i}=\tau_{i-1}\tau_{i}^{-1}$ ($\tau_{0}$ is the identity) and $\tau_{i}=\sigma_{i}^{-1}\sigma_{i-1}^{-1}\cdots\sigma_{1}^{-1}$. Using the package RAMP [2] for the software GAP [6], we can verify that $\mathcal{P}$ has simple Petrie polygons of length 12.
If $\mathcal{X}$ is the premaniplex in family 2 with 8 sides, the derived graph $\mathcal{X}^{\xi}$ from the voltage graph in Figure˜7 is a 5-polytope with 20736 octagonal faces, 82944 edges and 3456 vertices of degree 48. It has genus 29377.
To calculate the degree of the vertices we have to remember that a $4k$-step rotation (where $2k+1$ is the rank of $\mathcal{P}$) around a vertex of $\mathcal{X}^{\xi}$ corresponds to a 2-step rotation on the Petrie polygon of $\mathcal{P}$. So if $\ell$ is the length of the Petrie polygon containing the base flag of $\mathcal{P}$, the degree of the vertices of $\mathcal{X}^{\xi}$ is $2k\ell$ if $\ell$ is even, and $4k\ell$ if $\ell$ is odd. However, in odd ranks the length of the Petrie polygons of chiral polytopes cannotbe odd: this is because their existence would imply that there is a closed path of odd length in the maniplex, but each closed path must have the same number of flags in one orbit as in the other one.
If we were to prove that there are chiral $(2k+1)$-polytopes ($k\geq 2$) with simple Petrie polygons, ˜7.2 would imply the existence of polyhedra whose symmetry type graph is the premaniplex of family 2 with $4k$ sides. So we ask the following question
Question 1.
Are there chiral polytopes with simple Petrie polygons in any (odd) rank $n\geq 3$?
It might be enough to analyze the examples constructed by Pellicer in [19] or by Montero and Toledo in [18].
8 Interpretation via voltage operations